OpenStudy (anonymous):

sin^2*3a-sin^2*a = sin4a*sin2a Prove..

OpenStudy (anonymous):

two identities that CAN USE sin a + sin b = 2sin a+b/2 cos a-b/2 and sin a- sin b= 2 sin a-b/2 cos a+b/2 and third one 2sina cos a =sin2a

OpenStudy (anonymous):

yea.. but it on ^2

OpenStudy (anonymous):

den ur ALGEBRA is weak... :P PS- a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

im stuck.. give me another hint (;